OpenStudy (anonymous):

Directions solve for y slope-intercept from. 1) 5y-5=2x 2) 3x-4y=8 3) 2x+6y=-24

OpenStudy (gorv):

y slope intercept form y=mx+c

OpenStudy (gorv):

5y-5=2x 5y=2x+5 y=2x/5 +1